Cesar Rego Computers, a Mississippi chain of computer hardware and software retail outlets, supplies both educational and commercial customers with memory and storage devices. It currently faces the following ordering decision relating to purchase of disks:

Cesar Rego Computers, a Mississippi chain of compu
Answer #1

Given

Demand D = 36,500 disks

Setup cost S = $24

Holding cost H = $0.43

Purchase price P= $0.87

Discount price = $0.81

EOQ=\sqrt{2DS/H}

EOQ=\sqrt{(2*36500*24)/0.43}

EOQ = 2018.52 or 2019 units

Total cost at EOQ = Purchase cost + Holding cost + Ordering cost

= PD + (Q/2)H + (D/Q)S

= 0.87*36500 + (2019/2)0.43 + (36500/2019)24

=31755 + 434.085 + 433.88

= $32622.96

Now we will calulate Total cost at discounted price 0.81 and quantity Q = 5800

Total cost = 0.81*36500 + (5800/2)0.43 + (36500/5800)24

= 29565 + 1247 + 151.03

= 30963.03

Since the total cost with discounted volume is less then the total cost at EOQ, cesar rego should take the discount